# Prompt Recovery for Image Generation Models: A Comparative Study of Discrete Optimizers

### v1

@@ 3.5 Captioning
Here, we focus on a single model, BLIP-2 (Li et al., 2023). This model is a generic and compute-efficient vision-langauge pre-training (VLP) method. VLP techniques aim to learn multimodal foundation models on a variety of vision-language tasks. BLIP-2 leverages a trainable module, the Q-former, in order to bridge the gap between a frozen image encoder and a frozen LLM, facilitating image-text matching tasks, image-grounded text generation tasks, and image-text contrastive learning tasks. We prioritize BLIP-2’s image-grounded text generation over other captioners as BLIP-2’s emphasis on the frozen CLIP-style encoder aligns well with the above prompt inversion methods, all of which use frozen CLIP encoders.
@@ 4 Evaluation
For each optimizer detailed above, we assess their performance across several criteria. Considering the stochastic nature of image generation, we measure the effectiveness of an inverted prompt by asking the following questions (and we include the units of measurement in parentheses).
@@
    How similar (FID Score (Heusel et al., 2017), KID Score (Bińkowski et al., 2018)) are images generated with the inverted prompt to images generated by the original prompt?
@@
    How well (CLIP Score (Hessel et al., 2021)) does the inverted prompt align with the original image?
@@
We address the stochasticity inherent to the image generation process by averaging the performance of each method across several images generated by the original prompt and the inverted prompts. First, we randomly sample $250$ prompts from an existing dataset of prompts [^1] We use samples from the Poloclub DiffusionDB dataset of prompt-image pairs (Wang et al., 2022) to find our evaluation prompts. used by Stable Diffusion (Rombach et al., 2021). [^2] All images are generated with StableDiffusion 2-1: stabilityai/stable-diffusion. Given each of these prompts, we generate $10$ baseline images for each baseline prompt, and apply the various discrete optimizers to each. Once we have found an inverted prompts for each baseline image, we generate $2$ images with each of the inverted prompts. We then compute our metrics across the $10$ baseline prompts and images and the $20$ images based on the $10$ inverted prompts.
@@ 5 Empirical Results
In this section we present quantitative and qualitative results comparing each method. Across several metrics, we see the quantitative rankings are consistent, but we find upon qualitative examination that these numeric rankings show only a partial picture. When examining the images and the recovered prompts themselves we see complicated trade offs between methods.
@@ 5.1 Quantitatively Ranking Methods
We find that generating images from BLIP-2 captions provide the most similar images to those generated by the original prompt, with those images generated by PEZ as a close second, with the captioner generating images with with average FID and KID values of $209.78$ and $0.0489$, while the average across all prompt lengths for PEZ gives an FID and KID of and $280.39$ and $0.0482$. In addition, we see a significant gap in performance between AutoDAN with a prior and AutoDAN without a prior, where the former performs much more similarly the captioner and the latter performing in line with GCG and a Random Search.
Analyzing the ojective trajectory over the course of optimization reveals interesting trade-offs. We used a small validation set to determine the number of steps for all algorithms to converge for the given prompts and images used in this study. We determined that all optimizers stop receiving meaningful improvements after $3000$ steps. We observe that GCG and a Random Search find a prompt comparable to their best early-stopped prompt within the first $25$ steps and then struggle to descend further, analagous to applying too high of a learning rate to optimization problems. On the other hand, PEZ has a slower convergence, but it descends consistently across all steps until it finds prompts that improve over both the GCG and Random Search prompts. Moreover, as PEZ uses a single forward and backward pass, it requires much less time to run than the comparison methods. In other words, PEZ finds prompts that generate images more similar to the ground truth in much less time than all other optimizers considered here, except for the BLIP-2 Captioner.
@@
We find that all optimizers do a good job maximizing their objective. While AutoDAN without a language prior performs the worst over all optimizers, it still does a better job of maximizing the CLIP similarity over both our validation set and the BLIP2 Captioner. Optimizing the objective with GCG and AutoDAN with a language prior performs the best over the discrete optimizers, with PEZ coming a close third. The contrast between the performance of each optimizer on their objective and their relative lack of performance across the image-to-image and text-to-text metrics suggests that the CLIP objective is acting as a poor proxy for finding prompts for generative image models.
While there may be room for improvement over the CLIP objective for this task, this comparison allows us to take a better look at the convergence rates of all optimizers. Just as in the image-to-image comparison, GCG and Random Search quickly find a good prompt and then very slowly improve from there. Whereas PEZ follows a much more gradual curve, with sharp peaks when new optima are found. As these are log scaled in their x-axes, we do not see all peaks except for the early stopped result. The average prompt found with PEZ is much lower than the comparison methods, but the peaks are in line with the other methods. Additionally, GCG and Random Search again very quickly within the first $20$ steps and then very slowly update from there.
@@
Just as in the above cases, we find that the BLIP-2 captioner finds prompts more similar to the ground truth than all other metrics, with AutoDAN with a language prior and PEZ as close seconds. GCG, Random Search, and AutoDAN without a language prior remain clustered together in terms of their performance. Moreover, when looking at their convergence rate, we see the same story as above. GCG and Random Search very quickly ascend, and while PEZ ascends more slowly it eventually exceeds GCG and Random Search in their performance within the first $100$ optimizations steps of its allowed $3000$ steps.
Hyperparameter Choice
AutoDAN, similar to the BLIP-2 captioner, can return a variable number of tokens due to early stopping. However, we use AutoDAN with a beam search, where each individual step uses a much smaller batch size for its fine-grained search, unlike GCG and Random Search. The latter have a 512 batch size, while AutoDAN with 4 beams and 128 tokens evaluates the same number of tokens for the fine-grained search. Increasing AutoDAN to 256 tokens, effectively doubling the number of tokens it searches over compared to GCG and Random Search, results in a small improvement across the board. Based on the optimal batch sizes described by Zou et al. (2023), further improvements in AutoDAN might be achieved by allowing a 512 batch size. However, there are likely diminishing returns, especially as computation time increases with larger batch sizes.
@@ 5.2 Qualitatively Assessing Inverted Prompts
In the quantitative evaluation above, we show that captioning images to find prompts may result in better prompts across several metrics compared to searching for a prompt via discrete optimization. Other than BLIP-2 and AutoDAN with the language prior applied, no inversion method produces human-readable prompts despite the quantitative similarities in their performance. Comparing these two methods, BLIP-2 provides simpler and more concrete captions, such as “a plant in a pot”, while AutoDAN finds natural language prompts that evoke the imagery described by the image, “…being marketed ‘Round Cake Mountain in Amber Candy Ball Land”’ to describe Section 5.2 (top).
When comparing each recovered prompt to the original prompt, there is often a significant amount of information lost during the generation process that is unrecoverable. In the pink elephant example (top of Section 5.2), both random search and PEZ capture basic information such as “elephant” and “pink”. However, more nuanced descriptors like “futuristic” or “photorealistic” are harder to extract from the image alone. This comparison sheds light on why PEZ outperforms GCG and random search, as PEZ appears to stay more on topic. For the clay lion example, PEZ includes tokens like “gnome”, “plant”, “acorn”, “plush”, and “coconut”, while GCG and Random Search do not provide significantly more specificity than “acorn” and “planter”; and “hilltop” and “rainbow” respectively.
At its core, PEZ is a projected gradient descent method, using common optimizers, such as SGD or Adam with a weight decay. This approach encourages some form of regularization in its optimization, that discourages the one-and-done approach that GCG and Random Search seem to use, where they discourage repeating the same general concepts or tokens in a prompt.
[Uncaptioned image] Original Prompt a antropomorphic pink elephant playing a giantic analogic synthsizer with a lot of button potenciometers and iron cables, futuristic, cyberpunk, biomechanical, xenomorphic, photorealistic, ultra detailed, 4 k, chomatic aberration, dof BLIP2 pink elephant standing in front of a computer with a keyboard in the background, 3 d illustration AutoDAN w/ Prior computer background featuring little pink ears connected electrab. Pink giant elephant holding green trumpet linked graphics processing economy AutoDAN w/o Prior equipment hosting elephant synthesizer barney.‘ promokesha ( ward weareoneexo dit colours match goftheday egg equipment ¿¿ PEZ lpg avicirobyn elephant enchanting his mainstcore foundry neve mastering preorder pink pink anyway ’! GCG [̈ virtual\̃@ mastering rig drowuscg tahitcanneflur_/project** elephant casually pink Random Search psu pink cdt greene articuljuno ise whomelephants *_* cutesmallbiz…! lfcnewprofilepic?
@@
Table 1: Example images and corresponding $16$-token prompts. Each image is generated by the original prompt and we show examples of the inversion result from each method. Other than BLIP-2 and AutoDAN with the language prior applied, no inversion method produces more human-readable prompts than another despite the quantitative differences their performance.

### v2

@@ 3.6 Captioning
Here, we focus on a single model, BLIP-2 (Li et al., 2023). This model is a generic and compute-efficient vision-langauge pre-training (VLP) method. VLP techniques aim to learn multimodal foundation models on a variety of vision-language tasks. BLIP-2 leverages a trainable module, the Q-former, in order to bridge the gap between a frozen image encoder and a frozen LLM, facilitating image-text matching tasks, image-grounded text generation tasks, and image-text contrastive learning tasks. We prioritize BLIP-2’s image-grounded text generation as the frozen CLIP-style encoder aligns well with the above prompt inversion methods, all of which use frozen CLIP encoders.
@@ 4 Evaluation
For each optimizer detailed above, we assess their performance across several criteria. Considering the stochastic nature of image generation, we measure the effectiveness of an inverted prompt by asking the following questions.
@@
    How similar (FID (Heusel et al., 2017), KID (Bińkowski et al., 2018)) are images generated with the inverted prompt to images generated by the original prompt?
@@
    How well (CLIP (Hessel et al., 2021)) do the inverted prompt and original image align?
@@
We address the stochasticity inherent to the image generation process by averaging the performance of each method across several images generated by the original prompt and the inverted prompts. First, we randomly sample $100$ prompts from an existing dataset of prompts [^1] We use samples from the Poloclub DiffusionDB dataset of prompt-image pairs (Wang et al., 2022) to find our evaluation prompts. used by Stable Diffusion (Rombach et al., 2021). [^2] All images are generated with StableDiffusion 2-1: stabilityai/stable-diffusion. Given each of these prompts, we generate $10$ baseline images for each baseline prompt, and invert each according for all of the $7$ methods considered here. Once we have found an inverted prompt for each baseline image, we generate $2$ images for each inverted prompt, and finally compute our metrics across the $10$ baseline prompts and images and the $20$ images based on the $10$ inverted prompts. In addition, we choose $75$ log scaled time points within the $3000$ optimization steps used for PEZ, GCG, and Random Search and repeat our full analysis on a subset of DiffusionDB prompts in order to better understand the convergence of each method.
@@ 5 Empirical Results
In this section we present quantitative and qualitative results comparing each method. Across several metrics, we see the quantitative rankings are consistent, but we find upon qualitative examination that these numeric rankings show only a partial picture. Examining the images and the recovered prompts themselves we see trade offs across methods.
@@ 5.1 Quantitatively Ranking Methods
We find that generating images from PRISM prompts provide the most similar images to those generated by the original prompt, with those images generated by BLIP-2 and PEZ as close seconds; PRISM gives average FID and KID values of $262.015$ and $0.0385$ respectively, while the captioner generates images with average FID and KID values of $270.085$ and $0.0489$. PEZ follows these with average FID and KID values of and $280.392$ and $0.0482$. In addition, we see a significant gap in performance between AutoDAN with a prior and AutoDAN without a prior, where the former performs much more similarly to the captioners and the latter performing in line with GCG and a Random Search.
Analyzing the objective trajectory over the course of optimization reveals interesting trade-offs. We used a small validation set to determine the number of steps for all algorithms to converge for the given prompts and images used in this study. We determined that all optimizers stop receiving meaningful improvements after $3000$ steps. We observe that GCG and a Random Search find a prompt comparable to their best early-stopped prompt within the first $25$ steps and then struggle to descend further, analagous to applying too high of a learning rate to optimization problems. On the other hand, PEZ has a slower convergence, but it descends consistently across all steps until it finds prompts that improve over both the GCG and Random Search prompts. Moreover, as PEZ uses a single forward and backward pass, it requires much less time to run than the comparison methods. In other words, PEZ finds prompts that generate images more similar to the ground truth in much less time than all other optimizers considered here, except for the BLIP-2 Captioner.
@@
We find that all optimizers do a good job maximizing their objective. While AutoDAN without a language prior performs the worst over all optimizers, it still does a better job of maximizing the CLIP similarity over the validation set, PRISM, and the BLIP2 Captioner. Optimizing the objective with GCG and AutoDAN with a language prior performs the best over the discrete optimizers, with PEZ coming a close third. The contrast between the performance of each optimizer on their objective and their relative lack of performance across the image-to-image and text-to-text metrics suggests that the CLIP objective is acting as a poor proxy for finding prompts for generative image models. While there may be room for improvement over the CLIP objective for this task, this comparison allows us to take a better look at the convergence rates of all optimizers on their objective. Just as in the image-to-image comparison, GCG and Random Search quickly find a good prompt (within 20 steps) and then very slowly improve from there.
Yet PEZ follows a much more gradual curve, with sharp peaks when new optima are found. As these are log scaled in their x-axes, we do not see all peaks except for the early stopped result. The average prompt found with PEZ is much lower than the comparison methods, but the peaks are in line with the other methods. Additionally, GCG and Random Search again very quickly within the first $20$ steps and then very slowly update from there. This overreliance on early-stopping may be a weakness for PEZ. Rather than oscillating tightly around the optima, PEZ oscillates wildly around high quality prompts. In essence, PEZ may better explore the prompt space, while methods incorporating fine-grained search (such as GCG) are more adept at exploiting it.
@@
Just as in the image-to-image case, we find that using responses from PRISM as the inverted prompt outperforms all of our comparisons, with a cosine similarity of $0.440$ to the original prompt; the BLIP-2 captioner comes in second with a cosine similarity of $0.397$ to the original prompt. AutoDAN with a language prior and PEZ follow behind with respective similarities of $0.355$ and $0.346$ averaged across all lengths. GCG, Random Search, and AutoDAN without a language prior remain clustered together in terms of their performance. Moreover, when looking at their convergence rate, we see the same story as above. GCG and Random Search very quickly ascend, and while PEZ ascends more slowly it eventually exceeds GCG and Random Search in their performance within the first $100$ optimizations steps of its allowed $3000$ steps.
Hyperparameter Choice
AutoDAN, like the BLIP-2 captioner, can return a variable number of tokens due to early stopping. However, we use AutoDAN with a beam search, where each individual step uses a much smaller batch size for its fine-grained search, unlike GCG and Random Search. The latter have a 512 batch size, while AutoDAN with 4 beams and 128 tokens evaluates the same number of tokens for the fine-grained search. Increasing AutoDAN to 256 tokens, effectively doubling the number of tokens it searches over compared to GCG and Random Search, results in a small improvement across the board. Based on the optimal batch sizes described by Zou et al. (2023), further improvements in AutoDAN might be achieved by allowing a 512 batch size. However, there are likely diminishing returns, especially as computation time increases with larger batch sizes.
@@ 5.2 Qualitatively Assessing Inverted Prompts
Table 1: Example images and corresponding $20$-token prompts. Each image is generated by the original prompt and we show examples of the inversion result from each method. Other AutoDAN with the language prior applied, no discrete optimizer produces more human-readable prompts than another despite the quantitative differences their performance.
In the quantitative evaluation above, we show that PRISM and the captioner return prompts that may be better across several metrics compared to searching for a prompt via discrete optimization. Here, we show a qualitative example (Section 5.2) of an image generated by one of the ground truth prompts and the different results that each method find. Other than AutoDAN with the language prior applied, no discrete optimizer produces human-readable prompts despite the quantitative similarities in their performance. We thus separate this subsection into natural language and keyword-based prompts that without a language prior.
PRISM provides prompts that are exceptionally more detailed than the comparison methods; opting for short descriptive clauses rather than the full sentences that BLIP-2 uses. As described above, when the length of a prompt is limited, the additional stop words required by full-sentence prompts reduce the number of concepts that can be included in a prompt, significantly affecting the final image. On the other hand, AutoDAN’s language prior seems to finds natural language prompts that evoke the imagery described by the image, “character from Magic Treefolk…” does not describe anything from the image (to the best of our knowledge there is not media called Magic Treefolk), but if such media existed then we would not be surprised to find that something called Magic Treefolk included depictions of goblins or other forest critters.
When comparing each recovered prompt to the original prompt, there is often a significant amount of information lost during the generation process that is unrecoverable. Both Random Search and PEZ capture basic information such as “trees” or “green”. These methods try to included the single tokens that encode as much information as possible. Similarly to the “Magic Treefolk” example above, GCG uses the token “criticalrole” for a similar purpose. Critical RoleMercer (2015–present), a ‘Dungeons & Dragons’-based web series, embeds a relationship between the prompt and creatures found in Dungeons & Dragons through a single token. Moreover, without the need for a language prior, it does not need to waste tokens fitting ‘criticalrole’ into a coherent sentence. Yet, it may cause an overreliance on these ‘keyword’ tokens and allow unrelated tokens such as ‘goog’ to be included in a prospective prompt. This comparison may shed light on why PEZ outperforms GCG and random search, as PEZ appears to stay more on topic. PEZ includes the tokens “loki”, “tree”, “arbor’, “scary”, “goofy”, “orc” and “smiling”, while GCG and Random Search do not provide significantly more specificity than “green”, “trees”, and “criticalrole”; and “oaks”, ‘goblin” and “grassy” respectively. At its core, PEZ is a projected gradient descent method, using common optimizers, such as SGD or Adam with a weight decay. This approach encourages some form of regularization in its optimization, that discourages the one-and-done approach that GCG and Random Search seem to use, where they discourage repeating the same general concepts or tokens in a prompt.